OpenStudy (anonymous):

Choose two trinomials from the list below to factor. Using complete sentences, explain how to factor each one. Be sure that the final factorization (or "answer") is a part of your explanation. 2x2 + 13x + 15 2x2 + 27x + 13 5x2 − 2x − 7 10x2 − 7x − 3 8x2 − 11x + 3

jhonyy9 (jhonyy9):

8x2 -11x +3=0 x_1,_2=(11+/-sqrt(121-96))/16=(11+/- 5)16 = 6/16= 3/8 and 16/16=1 8x2 -11x +3= (x-3/8)(x-1)=(8x-3)(x-1)

jhonyy9 (jhonyy9):

5x2 -2x -7=0 x_1,_2 =(2+/-sqrt(4+140))/10 =(2+/-12)/10 = - 10/10 = -1 and 14/10 = 7/5 5x2 -2x -7=(x+1)(x-7/5)=(x+1)(5x-7)